OpenStudy (kitkat16):

The volume of a rectangular prism is b^3 + 8b^2 + 19b + 12 cubic units, and its height is b + 3 units. The area of the base of the rectangular prism is ?square units.

OpenStudy (kitkat16):

Can some one help me set this up? b^3 + 8b^2 + 19b + 12 = (b+1)(b+3)(b+4)

OpenStudy (kitkat16):

@iGreen are u busy?

OpenStudy (anonymous):

HELP PLEASE ON MY QUESTION I GREEN AND I DONT KNNOW HOW TO TAG

OpenStudy (igreen):

Just type in '@iGreen' on your question.

OpenStudy (igreen):

Okay, well we already have the length(b + 1) the width(b + 4) and the height(b + 3). Now we can plug this into the Surface Area formula. \(SA = 2(wh + lw + lh)\) \(SA = 2[(b + 4)(b + 3) + (b + 1)(b + 4) + (b + 1)(b + 3)]\) Multiply all 3 pairs of binomials together. \(SA = 2[(b^2 + 7b + 12) + (b^2 + 5b + 4) + (b^2 + 4b + 3)]\) Add like terms: \(SA = 2[(b^2 + b^2 + b^2) + (7b + 5b + 4b) + (12 + 4 + 3)]\) \(SA = 2[3b^2 + 16b + 19]\) Now multiply 2 to each term, can you do that?

OpenStudy (kitkat16):

im back sorry looking at it now thanks

OpenStudy (igreen):

That's fine, I just posted it.

OpenStudy (kitkat16):

so now I need to multiply 2 to each term? 2(3b^2+16b+19)

OpenStudy (kitkat16):

6b^2+32b+38

OpenStudy (igreen):

Yep. \(2(3b^2 + 16b + 19)\rightarrow (2 \times 3b^2) + (2 \times 16b) + (2 \times 19)\)

OpenStudy (igreen):

Yep, so your answer is \(SA = 6b^2 + 32b + 38\).

OpenStudy (kitkat16):

I wish but that wasn't one of my answers. Ive been going crazy since yesterday.

OpenStudy (igreen):

What..what are your answer choices?

OpenStudy (kitkat16):

b^2+5b+4 (b^2+5b+4+24)/(b+3) (b^2-5b+4-24)/(b+3) (-b^2+11b-52+168)/(b+3)

OpenStudy (igreen):

Ohhhhhh.....

OpenStudy (kitkat16):

see insane!!!!!

OpenStudy (igreen):

Hold on, let me test something out.

OpenStudy (kitkat16):

ok thanks

OpenStudy (igreen):

Okay, I have no idea..lol. Sorry..

OpenStudy (kitkat16):

Its OK thanks for trying I really appreciate it!!!

OpenStudy (igreen):

@Supreme_Kurt @ganeshie8 @campbell_st @iambatman @perl

OpenStudy (anonymous):

Actually, @KitKat16 You've already done the hard part of this blasted question XD You've already factored the thing, and you know which factor is the height... so.... We know that the length and width of the rectangular base are (b+1) and (b+4) ...right?

OpenStudy (kitkat16):

yes

OpenStudy (kitkat16):

@Supreme_Kurt all my questions are insane ask @iGreen lol

OpenStudy (anonymous):

So, you're asked for the AREA of the rectangular base, and you know the length and width... Can't get any smoother than that ;)

OpenStudy (anonymous):

One tag's enough KitKat, just call me Kurt ^

OpenStudy (kitkat16):

(b+1)b+3)(b+4)

OpenStudy (kitkat16):

ok

OpenStudy (igreen):

You got the length, width, and height.

OpenStudy (anonymous):

Why'd you include b+3? That gives you the volume lol The length and width of the rectangular base are just b+1 and b+4 . So...

OpenStudy (kitkat16):

b^2+5b+4

OpenStudy (anonymous):

Exactly ;) Nice job.

OpenStudy (kitkat16):

omg it was that easy. Thanks guys you are awesome! Thanks for always being there Green

OpenStudy (igreen):

Hehe, no problem :)